6. tx - 2x + tx = 0, x(0) = 0.

Use the Laplace transform to solve initial value problems
Answer #1

Solution:

Notice first that taking t = 0 into the given equation tx - 2.r + tx = 0 we get 0 - 2x(0) + 0 = 0. it follows that xO) =

Taking the inverse Laplace transform we get z(t) = 2-1 e-{co2+1;} Csint - tcost) (s2 +

where C + 0 is an arbitrary constant.

Thank you